OpenStudy (anonymous):

Help me identify the coordinates of the center of this ellipse. (x-2)^2/36+(y+1)^2/25=1

OpenStudy (jdoe0001):

\(\bf\cfrac{(x-h)^2}{a^2}+\cfrac{(y-k)^2}{b^2}=1\\ \textit{center is at coordinates }(h,k)\)

OpenStudy (anonymous):

\[\frac{(x-\color{red}h)^2}{a^2}+\frac{(y-\color{blue}k)^2}{b^2}=1\] \ \[\frac{(x-\color{red}2)^2}{a^2}+\frac{(y-\color{blue}{(-1}))^2}{b^2}=1\]

OpenStudy (anonymous):

center is \((\color{red}h,\color{blue}k)\)

OpenStudy (anonymous):

therefore the coordinates are (2,1) , correct?

OpenStudy (anonymous):

sorry (-2,1)

OpenStudy (jdoe0001):

\(\bf \cfrac{(x-\large{\color{red}2})^2}{a^2}+\cfrac{(y-\color{blue}{(\large{-1}}))^2}{b^2}=1\) look closely

OpenStudy (jdoe0001):

\(\bf \cfrac{(x-h)^2}{a^2}+\cfrac{(y-k)^2}{b^2}=1\)

OpenStudy (anonymous):

ok so would it be (2,-1) ?

OpenStudy (jdoe0001):

yeap, the " - " in the equation, isn't part of the "h" or the "k", so yes, 2, -1

OpenStudy (anonymous):

thanks!

OpenStudy (jdoe0001):

yw